OpenStudy (anonymous):

what is the 15th term of the geometric sequence -√5, √10, -2√10

OpenStudy (jdoe0001):

do you know what a geometric sequence is?

OpenStudy (anonymous):

yeah its a sequence with a common ratio

OpenStudy (anonymous):

I just suck at radicals :(

OpenStudy (jdoe0001):

got any ideas what the common ratio might be?

OpenStudy (anonymous):

I got an equation an=a1 x r ^(n-1)

OpenStudy (anonymous):

a1= first term, r= common ratio and n=term

OpenStudy (jdoe0001):

anyhow the difference between "a term" and the "next term" is a multiplier so if you just divide the "next term" by the "a term" then your quotient would be the multiplier or "common ratio" so let us use the 1st and 2nd term to get it \(\bf \cfrac{\sqrt{10}}{-\sqrt{5}}\implies \cfrac{\sqrt{2\cdot 5}}{-\sqrt{5}}\implies \cfrac{\sqrt{2}\cdot \cancel{\sqrt{5}}}{-\cancel{\sqrt{5}}}\) can you see the common ratio ?

OpenStudy (anonymous):

is it √2 ?

OpenStudy (jdoe0001):

well.... notice the - at the bottom, so it's negative

OpenStudy (jdoe0001):

so \(\Large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1} \\ \quad \\ a_1=\textit{first term in the sequence}\qquad r=\textit{common ratio} \\ \quad \\ \textit{so the 15th term will be }\implies \Large a_{\color{red}{ 15}}=a_1\cdot r^{{\color{red}{ 15}}-1}\)

OpenStudy (jdoe0001):

\(\Large \bf a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ n}}-1} \\ \quad \\ a_1=-\sqrt{5}\qquad r=-\sqrt{2} \\ \quad \\ \textit{so the 15th term will be }\implies \Large a_{\color{red}{ 15}}=-\sqrt{5}\cdot (-\sqrt{2})^{{\color{red}{ 15}}-1} \)

OpenStudy (anonymous):

so -√5 X -√2^14

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

but I don't know how to solve (-√2)^14

OpenStudy (jdoe0001):

hmm \(\bf a_{\color{red}{ 15}}=-\sqrt{5}\cdot (-\sqrt{2})^{{\color{red}{ 15}}-1} \\ \quad \\ (-\sqrt{2})^{14}\implies (-1\cdot \sqrt{2})^{14}\implies (-1)^{14}\cdot (\sqrt{2})^{14}\implies 1\cdot \sqrt{2^{14}}\)

OpenStudy (anonymous):

the answer choices are 1) -128√5 2) 128 √10 3) -16384 √5 4) 16384 √10

OpenStudy (jdoe0001):

so what would you get from say \(\bf (-\sqrt{2})^{14}\implies (-1\cdot \sqrt{2})^{14}\implies (-1)^{14}\cdot (\sqrt{2})^{14}\implies 1\cdot \sqrt{2^{14}}\) ?

OpenStudy (anonymous):

you would get √16384 and so how would you do how would you do √16384 x -√5

OpenStudy (jdoe0001):

well.... you wouldn't

OpenStudy (jdoe0001):

but you can get square root of 16384

OpenStudy (anonymous):

so -128 √5 is the answer right

OpenStudy (jdoe0001):

\(\bf (-\sqrt{2})^{14}\implies (-1\cdot \sqrt{2})^{14}\implies (-1)^{14}\cdot (\sqrt{2})^{14}\implies 1\cdot \sqrt{2^{14}} \\ \quad \\ \sqrt{2^{14}}\implies \sqrt{(2^7)^2}\implies \Large \sqrt[{\color{red}{ 2}}]{(2^7)^{\color{red}{ 2}}}\implies 2^7\)

OpenStudy (anonymous):

2^7 = 128 so 128 x -√5 = -128 √5 right ?

OpenStudy (anonymous):

thanks for your help